OpenStudy (anonymous):

am i correct in thinking the only solution to cos(x) - 1 + ((x^2)/2) = 0 is x = 0 ?

OpenStudy (anonymous):

only real*

OpenStudy (datanewb):

Maybe take the derivative to get a better idea?

OpenStudy (anonymous):

well here is my argument, i dont know whether it's totally valid: consider f(x) = sinx - x f'(x) = cosx -1 =< 0 therefore f(x) is decreasing, since f(0) = 0 we have sinx - x <0 for x<0 ie: sinx< x now consider sinu < u sin^2 (u) < u^2 -2sin^2 (u) < -2u^2 1- 2sin^2 (u) < 1 - 2u^2 cos(2u) < 1 - 2u^2 let u = x/2 now cos(x) < 1 - (x^2)/2 for 0<x since cos(x) and 1 - (x^2)/2 are even we can say the same for x<0 therefore the only real solution for x is x = 0

OpenStudy (datanewb):

Your logic seems correct to me. I was thinking something like: \[f(x) =cos(x) - 1 + ((x^2)/2)\] \[f(0) = 0 \\ f'(x) = -sin(x) + x \\ f'(0) = 0 \\ f''(x) = -cos(x) + 1 \\cos(x) \leq 1 \Rightarrow f''(x) \geq 0 \Rightarrow f'(x) \geq 0 \Rightarrow f(x) \geq 0 \] And thus, 0 is the only solution to the original equation. Although, I think maybe your way is just as good if not better.